### v1

# LLMs in the SOC: An Empirical Study of Human-AI Collaboration in Security Operations Centres

Abstract
The integration of Large Language Models (LLMs) into Security Operations Centres (SOCs) presents a transformative, yet still evolving, opportunity to reduce analyst workload through human-AI collaboration. However, their real-world application in SOCs remains underexplored. To address this gap, we present a longitudinal study of 3,090 analyst queries from 45 SOC analysts over 10 months. Our analysis reveals that analysts use LLMs as on-demand aids for sensemaking and context-building, rather than for making high-stakes determinations, preserving analyst decision authority. The majority of queries are related to interpreting low-level telemetry (e.g., commands) and refining technical communication through short (1-3 turn) interactions. Notably, 93% of queries align with established cybersecurity competencies (NICE Framework), underscoring the relevance of LLM use for SOC-related tasks. Despite variations in tasks and engagement, usage trends indicate a shift from occasional exploration to routine integration, with growing adoption and sustained use among a subset of analysts. We find that LLMs function as flexible, on-demand cognitive aids that augment, rather than replace, SOC expertise. Our study provides actionable guidance for designing context-aware, human-centred AI assistance in security operations, highlighting the need for further in-the-wild research on real-world analyst-LLM collaboration, challenges, and impacts.
1 Introduction
Figure 1: SOC workflow, focus of study and our insights.
SOCs handle diverse functions, from real-time detection and incident response to continuous improvement [1, 2, 3]. SOC analysts play a critical role in this by interpreting data and making informed decisions. They typically follow a multi-stage process, beginning with quick assessments of alert relevance [4], then checking for related signals, gathering contextual telemetry to build situational awareness [3], and evaluating evidence of compromise to make escalation decisions [5]. This reasoning workflow is supported by tooling like SIEM, SOAR, and XDR [3, 6].
Recently, LLMs have emerged as promising tools to support such SOC workflows, with applications ranging from threat intelligence to triage automation [7, 8, 9, 10]. While prior research explores human-LLM interaction in various domains and lab studies [11, 12, 13, 14], it does not capture how SOC analysts utilise LLMs in their daily workflows. Existing work, such as Copilot Guided Response (CGR) [7], show system-level performance, not individual analysts’ usage patterns.
Gaining insights into how SOC analysts utilise LLMs in practice, including the types of questions they ask and the tasks they use them for, is essential for evaluating the practical role of LLMs in operational cybersecurity (Figure 1). This study addresses that gap by analysing 3,090 queries submitted to an internally deployed instance of GPT-4 over 10 months (May 2023 to March 2024) by 45 SOC analysts at eSentire Inc., a cybersecurity firm specialising in 24/7 managed detection and response. To guide this investigation, we focus on the following research question:
How do SOC analysts utilise LLMs in their daily workflows, including the specific tasks they apply them to, how these uses align with established cybersecurity frameworks (e.g., NICE), and the conversational patterns that characterise their interactions?
The answers to this question reveal SOC analysts’ task priorities and engagement patterns with LLMs, informing the design of more effective LLM-based assistants as human-AI collaboration becomes increasingly common in SOC environments. We address this question through a multi-level analysis of analyst-LLM interactions across queries, tasks, and conversations.
  • Query-level: We examine what analysts ask and how they phrase their requests, revealing their cognitive needs, the daily tasks augmented by LLMs, and alignment with frameworks such as NICE.
  • Conversation-level: We analyse sequences of queries within investigative sessions to uncover how analysts structure multi-turn interactions in their daily workflows. This reveals reasoning patterns and functional roles that LLMs assume in conversations.
  • Analyst-level: We investigate the frequency and manner in which different analysts use LLMs. This variation reveals diverse strategies for integrating LLMs and highlights how they become embedded in the daily workflows of different analysts.
We began with descriptive analysis to map overall query activity, including how frequently analysts interacted with the LLM. Then, we conducted a thematic analysis [15] to identify patterns in query phrasing, topics, and tasks. Since our research objective was to characterise analyst behaviour, not evaluate LLM performance, we focused our thematic analysis on analyst queries and used LLM responses only to contextualise intent. We grouped successive analyst queries on the same topic into single conversational units, based on topic continuity. Finally, we synthesised insights across Query, Conversation, and Analyst levels to characterise analyst-LLM interactions. Although our dataset covers a single enterprise SOC using one type of text-based LLM, it provides a valuable longitudinal view of real-world LLM use, highlighting key usage patterns and tasks. As more studies emerge, both quantitative (e.g., [7]) and qualitative (e.g., ours, [16, 5]), research and design communities will be better informed to design targeted LLM-based SOC tools.
Contributions
We make the following contributions:
  • We present the first in-depth, real-world analysis of LLM adoption in a SOC, spanning 3,090 queries from 45 analysts over 10 months. Our analysis highlights analyst-level usage patterns across task types and individual differences.
  • We develop a structured coding framework that facilitates thematic analysis of analyst-LLM interactions by classifying LLM queries according to task type, subject, and interaction pattern. We apply this framework to develop a deeper understanding of how LLMs assist investigative, explanatory, and communicative functions within SOC workflows.
  • We provide empirical insights into analyst behaviour and LLM utility in SOCs. We identify three core use cases that characterise how SOC analysts engage with LLMs: (i) interpreting raw telemetry, (ii) generating and refining task-related communication content, and (iii) obtaining rapid, on-demand assistance through concise interactions. Collectively, our analysis reveals that LLMs are becoming routine, competency-aligned aides rather than decision-makers: 93% of queries map to NICE tasks, usage has shifted from ad-hoc exploration to sustained integration, and analysts retain final judgement while leveraging LLMs for rapid sense-making. These findings guide the design of context-aware, human-centred assistants to enhance human-AI collaboration in SOCs.
  • We offer design recommendations for AI-augmented SOCs: (1) embed analyst- and context-adaptive LLM explanations of technical artifacts to speed triage and support evidence-based decisions; (2) integrate LLM assistants for microtasks to provide in-workflow support and minimise context switching; and (3) surface evidence over recommendations for investigative tasks, preserving analyst autonomy and context.
2 Related Works
2.1 LLMs for Cybersecurity
Prior to LLMs, SOCs already leveraged AI across SIEM correlation, IDS/IPS, and related tooling [17]. Recent works highlights the expanding role of LLMs in cybersecurity workflows. For example, Copilot Guided Response (CGR) aids triage and remediation by generating recommendations from historical threats [7], while tools like LocalIntel [18] extract structured threat intelligence from unstructured sources to accelerate threat contextualisation. LLMs also support vulnerability scanning [19, 20, 21], anomaly detection [22, 23], intrusion detection[24], phishing simulation [25, 26, 27], threat intelligence summarisation [28] and privacy-preserving analysis [29], leveraging their generalisation and pattern recognition capabilities. However, most of these systems are evaluated in controlled environments, with a focus on performance metrics such as speed and accuracy of decision-making (e.g. [7]). In contrast, our study provides the first empirical analysis of LLM usage in a live SOC, examining how analysts actually engage with the LLMs, what help they seek, how interactions unfold, and how adoption evolves over time, offering a unique, human-centred view into analyst-LLM interactions in practice.
2.2 LLM Roles in Decision Support
LLMs play diverse roles in human-AI decision-making, shaped by system autonomy, user needs, and interaction structure [30, 31, 32, 33, 34]. LLMs span roles from assisted intelligence (retrieval and summarisation) to more collaborative forms like augmented and cooperative intelligence, which aid reasoning and joint problem structuring [30, 33]. HCI research highlights the importance of aligning AI support with decision phases and initiatives, whether recommending actions, making predictions, proposing alternatives, or surfacing evidence [35, 32, 31, 33]. While AI suggestions can speed decisions, they may also increase over-reliance [36, 37]. Cognitive forcing techniques (e.g., delaying AI output) can reduce over-reliance but risk under-reliance among experienced users [36, 38]. Evaluative AI offers a middle ground by surfacing pros and cons without prescribing decisions [39], supporting autonomy in machine-in-the-loop systems [40, 41, 42]. Our analysis examines how LLMs operate in practice and the types of decision support that analysts seek.
2.3 Interaction Patterns and Analyst Strategies
Despite the growing capabilities of models, interaction patterns between humans and LLMs remain underexplored. Studies show that interaction framing and modality shape user expectations, trust, and behaviour [13, 43, 12, 11, 44, 45, 27]. Recent work highlights that prompt structure shapes explainability and collaboration [13], dialogue in complex tasks involves verbal and nonverbal cues [44], and human–AI interaction spans planning to testing across varying agent roles [11]. These studies underscore that the effective integration of LLMs for decision-making hinges on their ability to flexibly assume different roles shaped by user needs, interaction design, and evolving task context. Our study addresses this gap by analysing thousands of analyst-LLM interactions in an operational SOC, uncovering concrete interaction patterns, evolving strategies, and the situated use of LLMs as cognitive aids in high-stakes, time-sensitive environments.
3 Method
To analyse and understand SOC analysts’ interactions with LLMs, we adopted a six-phase methodology, as illustrated in Figure 2, which we describe in detail below.
Phase 1: Exploratory Analysis of Interactions. In this phase, we performed a statistical analysis of analyst-LLM interactions to characterise overall usage patterns, such as the total number of queries, engagement frequency, and query length distributions, providing a quantitative foundation for understanding analyst behaviours before moving to deeper thematic and qualitative analysis.
Phase 2: Data Familiarisation for Question Coding & Conversation Tagging (Partial Dataset). This phase involved three members of the research team reviewing the question-LLM response pairs multiple times to develop a shared understanding of the analyst queries as well as identify and tag conversations for thematic analysis. Since the data consisted of individual analyst-LLM interactions, the team first needed to group these into coherent conversations.
Question Coding. This step focused on developing a coding scheme to analyse the nature of the analysts’ queries. The three researchers first familiarised themselves with the data by reviewing a range of question-LLM response pairs. Through discussion, the team agreed on three key attributes to enable a meaningful decomposition of the queries:
  • Query Pattern: The structural form or phrasing of the analyst’s input, abstracted to remove specific details (e.g., filenames, commands, URLs). For example, a pattern such as “What is this $\ldots$] doing”, captures a common syntactic form regardless of the specific artifact used. This abstraction facilitates grouping similar queries and supports broader pattern analysis.
  • Query Subject: The primary topic or focus of the question, indicating the subject matter the analyst is asking about (e.g., a command, malware, or file).
  • Task: The underlying intent behind the query, categorised into functional areas such as Command Understanding or Summarising LLM Output.
To ensure consistency and reliability in applying this coding scheme, a sample of 20 question-LLM response pairs was independently coded by all three researchers. This was followed by a calibration session where the same 20 pairs were jointly reviewed and coded to to reconcile differences and refine definitions. Once alignment was achieved, the researchers independently coded a further 310 (10%) question-response pairs to test their collective understanding of the attributes and coding scheme. The reconciliation sessions helped resolve any disagreements and update the codebook. IRR, measured using Fleiss’ Kappa across the three dimensions, showed strong agreement: 0.90 for Query Pattern, 0.82 for Query Subject, and 0.79 for Task, indicating substantial to near-perfect agreement.
Conversation tagging. The next step involved grouping successive analyst queries that relate to the same topic or investigative thread and treating them as a single conversational unit by considering (a) overlapping security artifacts (e.g., commands, email rules), (b) continuity in investigative intent (e.g., both steps probing the same log entry), and (c) occurring within a 30-minute window (used only as a guideline rather than a strict cutoff). Each conversation was assigned a unique identifier (e.g., C1, C2). To ensure consistency and reliability in tagging, the three researchers independently labelled a random sample of 45 analyst questions, comprising 142 question–response pairs, the met to reconcile and refine their understanding of the tagging process. Once consistent tagging practices were established, an additional 240 instances (13%) were tagged as a further verification step. Fleiss’ Kappa was used to calculate IRR, resulting in a score of 0.75, indicating substantial agreement.
Figure 2: The six-phased approach to analyse and understand SOC analysts’ interactions with LLMs.
Phase 3: Remaining Dataset Question Coding & Conversation Tagging. In this phase, given acceptable agreement in Phase 2, the remaining question-LLM response pairs were divided among the three researchers, with each coding approximately 900 pairs independently. After initial coding, each researcher’s annotations were cross-verified by another team member to ensure consistency and accuracy; discrepancies were resolved through discussions. To further enhance the reliability of the coding process, one researcher conducted a comprehensive review of all the coded data. Discrepancies were minimal, observed only in 4% of the cases and resolved collaboratively through discussions.
Phase 4: Query Level Thematic Analysis. One researcher consolidated all analyst question-LLM response pairs into a dataset and conducted query level thematic analysis [15], as outlined in Figure 2. This phase built on earlier manual coding (Phase 3) and introduced a hybrid approach that combined qualitative coding with data-driven clustering and thematic interpretation (Phase 5) to surface higher-level insights. We applied semantic clustering to the coded data to support theme development. Sentence-BERT embeddings (stsb-roberta-large) and agglomerative clustering with cosine similarity [46] were used to group over 1000 distinct query patterns into semantically similar codes, supporting pattern abstraction. All clusters were manually reviewed to ensure semantic coherence. Similar patterns such as explain [x] and explain this [x] were grouped. Themes were then developed by merging related clusters and assigning meaningful, high-level labels spanning across all attributes; e.g., command and multiple commands formed a broader command-related subject theme.
To ensure the robustness of these themes, we conducted a triangulation process [47] in which two additional researchers, who were not involved in the clustering, independently reviewed and refined the themes. This hybrid approach, combining manual coding, computational clustering, and human interpretation, ensured both methodological rigour and interpretive depth [48, 49].
Phase 5: Conversation and Analyst Level Thematic Analysis. In the final phase, we conducted an in-depth analysis using the final annotated dataset from Phase 4. This involved both conversation-level and analyst-level analysis to uncover patterns at both levels. We first analysed the structure, frequency and patterns of conversations, identifying common interaction styles and task transitions. Analysts were then clustered based on usage volume, revealing distinct engagement profiles and thematic preferences. Finally, we explored behavioural and cognitive patterns, including low-engagement users, to understand diverse styles of LLM integration and potential disengagement plans.
4 Results
Figure 3: (Left) Query volumes vary across analysts, with heavy concentration among a few users. (Right) Overall, there is growing integration into workflows, but mostly driven by a subset of analysts (March 2024 has only 7 days of data).
This section presents findings from our analysis of LLM usage. We start with an overview of SOC analysts, model information, and the dataset before diving into exploratory analysis. We then present the findings from a three-level thematic analysis covering query, conversation, and analyst levels. The dataset comprised 3,122 analyst-submitted queries to GPT-4, collected between May 18, 2023, and March 7, 2024. After removing 32 queries where the LLM could not provide a meaningful response (e.g., due to missing input or unsupported tasks), 3,090 valid queries and 532 distinct conversations remained for analysis.
Analyst and Model Information
The 45 SOC analysts comprised: 1) SOC Analyst I: 52%; 2) Senior SOC Analyst: 17%; 3) Threat Analyst: 14%; 4) SOC Analyst II: 8%; 5) SOC Incident Handler: 2%; 6) Associate Analysts: 7%. All participants, except Associate Analysts, had 3–5 years of experience and worked with a proprietary SOC stack integrating alerts, logs, and threat intelligence. They used GPT-4 (GPT-4-0613) was used during live investigations without fine-tuning or access to the internet or SOC systems. Although other models were available, our analysis focuses solely on GPT-4. Analysts were free to use the LLM as they saw fit, but were instructed not to share sensitive or identifying data.
Figure 3 (left) summarises analyst engagement with the LLM over 10 months. Usage was uneven, with a small number of analysts generating the majority of queries. Given the skewed usage, we clustered the analysts based on their usage before analysing their long-term reliance on the LLM. We applied the Fisher-Jenks natural breaks algorithm [^1] https://en.wikipedia.org/wiki/Jenks_natural_breaks_optimization to total query counts per analyst. This generated three distinct clusters: a large group of low-usage analysts (Cluster 0, $n=34$), a smaller group of moderate users (Cluster 1, $n=10$), and a single power user (Cluster 2, $n=1$; most active analyst). Figure 3 (right) shows that daily usage increased steadily over time, driven by a subset of highly active analysts, suggesting deeper integration of the LLM into daily workflows. The number of active analysts also grew month on month, suggesting sustained interest and continuous adoption of the LLM.
4.1 Exploratory Analysis
While a single power user submitted approximately 600 queries (17% of the total dataset), usage was not limited to a few individuals. By early 2024, over 45 analysts had interacted with the system, and daily query volumes increased from fewer than 10 queries per day in May-August 2023 to over 30 queries per day in January and February 2024. We analysed the monthly usage patterns and observed that the number of active analysts grew from 4 in May 2023 to 13 in June, reaching 18 by July and steadily increasing to a peak of 25 in February 2024. In just the first few days of March 2024, 16 analysts had already submitted queries, indicating continued high engagement.
The monthly increases include the proportion of returning analysts and new analysts rising each month (see Figure 8 in the Appendix). By early 2024, the majority of LLM interactions came from returning analysts, indicating ongoing integration of the tool into daily workflows. Most active analysts utilise the LLM in tightly clustered bursts, with a median gap of 1-2 hours between visits, suggesting its use as an on-demand aid between investigative steps.
The average query length was 25 words (standard deviation: 70), but this distribution was skewed by a few very long queries. Some analysts submitted highly verbose queries, exceeding 1000 words (due to code or script segments), while others consistently relied on concise prompts of fewer than 20 words. In contrast, LLM responses were more uniformly verbose, with a mean of 161 words (standard deviation: 97) and a relatively narrow interquartile range.
4.2 Query Level Analysis
This section discusses the thematic analysis results, which involved coding the individual queries. Note that each conversation could involve multi-step interactions. The thematic analysis was conducted on individual queries, rather than entire conversations, to avoid missing the granular insights these provide over a holistic understanding of conversations. Section 4.3 presents conversation-level analysis. A complete list of themes can be found in the appendix.
The findings from the exploratory analysis reveal an interesting landscape of LLM use in our dataset. Despite wide variation in query volume, engagement style, and query lengths, usage trends point to a transition from periodic exploration to routine integration, with increased adoption over time and sustained interaction by a subset of analysts stronger increase among 11/45. These observations offer grounded insights into how LLMs are appropriated in practice, setting the stage for the next phase of our study: a qualitative analysis of the tasks and interaction strategies underpinning analyst-LLM collaboration.
Analysts varied widely in usage patterns, with some submitting over 600 queries and others fewer than 5. To assess the robustness of our results to outlier effects following thematic analysis, we recalculated the proportion of queries assigned to each task theme after excluding (i) the most active analyst, (ii) the top 10 most active analysts, and (iii) both the most active and the 20 least active analysts. Across all scenarios, the rank order of the top ten and the top two themes remained unchanged.
Figure 4: Summary of Task Themes (top 10)
4.2.1 SOC Analyst Task Types
Task themes reflect the actual work that analysts aim to accomplish when interacting with an LLM. We find that analysts used LLMs for a range of tasks central to day-to-day SOC operations [1, 3]. These include interpreting and analysing commands; editing and improving the clarity of incident descriptions, documentation, and alert text; and understanding or troubleshooting code, scripts, and regular expressions. Other common uses included analysing files and malware, exploring tool and system behaviours, constructing queries for data lookup, and clarifying email rules or suspicious artifacts, highlighting the breadth of tasks analysts juggle in operational contexts. Below, we detail three of the most common task themes (see Figure 4 for the top 10) illustrating how LLMs are integrated in operational workflows (see Figure 9 in the appendix to observe how task adoption changed over time).
Command Understanding, Analysis & Generation
This was the most frequent task theme (31% of queries). Queries typically involved analysts submitting command-line instructions and requesting clarification on functionality, syntax, or behavioural purpose. Common prompts such as “what does [x] do” or “explain this [x]” ([x] is the actual command(s)) suggest investigative or verification-oriented goals, with the LLM supporting the analyst through interpretive scaffolding. While we did not perform a thorough thematic coding of the LLM responses, we did analyse the responses it provided for the command-related questions to understand why this theme was likely prominent.
The LLM often began with an overview of the command, explaining the components and their relationships, inferring the intent behind complex operations, and highlighting suspicious behaviours. For example:
“In summary, the [anonymised command] is trying to login as [anonymised user] without password and $\ldots$
Many responses also included parameter breakdowns, example outcomes, risk implications, and recommended follow-up actions. Based on their sustained usage, we hypothesise that the LLM helped by decomposing command components and synthesising likely intent and behaviour, , the LLM has shown the potential of enhancing the analysts’ situational understanding. This aligns with prior work (e.g., [16]), which finds that SOC practitioners prefer explanations that contextualise evidence and inform action.
Text Generation & Editing
This theme, comprising 22% of queries, included requests to rephrase incident descriptions or alert summaries or correct grammar. Analysts frequently used queries such as “make this better”, “can you correct this”, or “rephrase”. This theme underscores the increasing expectation that LLMs will assist SOC analysts in producing clear, coherent written documentation for SOC systems and peer or client communication.
Code, Script, & Regex Analysis
Representing 11% of queries, this theme included submissions of code blocks (e.g., Python, PowerShell, JavaScript) and regex expressions. Analysts requested help understanding their behaviour, decoding strings, or generating new logic. Example queries included “what does this script do”, or “write a regex to detect [x]”. This theme suggests that analysts and the LLM engaged in complex reasoning around syntax, logic, and security behaviour of scripts and code. Regex generation was also notable, typically used in filtering logs.
Other Technical Tasks
In addition to the top three themes, analysts used LLMs for a wide range of other tasks including Tool/ Software/ System Understanding; Attack, Security & Threat Intelligence; Access Control & Authentication; File & Malware Analysis; Network & Traffic Analysis; and Query, Search & Data Processing. These tasks reflect specialised needs where LLMs provide targeted support. Notably, LLM Chat interactions suggest a growing social relationship between analysts and the assistant; five analysts explicitly thanked the LLM.
The results reveal a growing reliance on LLMs across multiple functions: from interpreting low-level artifacts, to generating or refining work products such as incident notes, queries, and email, as well as conducting on-demand look-ups, including definitions. Findings suggest that LLMs are emerging as flexible cognitive aids (e.g., as explainer or interpreter, drafting assistant, coding helper, and on-demand reference tool) rather than just static chatbots. The diversity of tasks calls for a unified integration.
4.2.2 SOC Analyst Query Subjects
We encoded query subjects to preserve anonymity (to not disclose proprietary information), resulting in a strong one-to-one mapping between subjects and task themes. For example, the task Command Understanding aligns with the subject command.
Across the 3,090 coded queries, subject matter clustered into a few dominant types: command-related artifacts (e.g., powershell and curl commands) led by a wide margin ($\approx$30%), followed by requests about phrasing or formatting snippets of written content ($\approx$17%) and code or scripting issues such as regex or PowerShell ($\approx$9%). Tool-specific questions ($\approx$7%) and file- or directory-path look-ups ($\approx$5%) rounded out the top five. The remainder was dispersed across LLM meta-chat, malware or threat look-ups, MDR/EDR query tuning, alert narration, network indicators, email-rule checks, and assorted ‘how-to’ prompts.
Telemetry tells us what was observed; analyst queries reveal which aspects they seek to interpret using LLMs. While security teams collect rich telemetry via tools like SIEMs and EDRs, LLMs are increasingly used to interpret parts of that raw input, to potentially transform the raw input into actionable insights or contextual understanding. Analysts appear to have relied on LLMs to interpret telemetry-derived artifacts (e.g., commands, log fragments), to likely improve situational awareness and support timely decision-making.
The MITRE ATT&CK data sources [^2] https://attack.mitre.org/datasources/ describe telemetry collected by sensors (e.g., process creation, file access, authentication logs). Most analyst queries focused on interpreting artifacts that these sensors collect, such as commands, scripts, OS processes or file paths. While subject themes do not directly map to MITRE ATT&CK data sources, they reflect reasoning tasks shaped by this telemetry. For instance, when an analyst asked the LLM to explain a PowerShell command launching a scheduled task, that artifact maps to ‘Command Execution’.
Figure 5: Summary of Pattern Themes (top 10)
4.2.3 SOC Analyst Query Patterns
Figure 5 shows how analysts linguistically framed their prompts. The most frequent pattern, Functional Understanding (21%), included queries like “what does [x] do”. These were used to interpret unfamiliar commands, scripts, or processes. Text Processing and Editing (16%) reflected efforts to refine reports or alerts (“reword this”), while Definition and Explanation requests (15% and 14%) were used to close knowledge gaps. Less common but operationally meaningful patterns included Content Generation (e.g., writing rules or code) and Validation/Correction (e.g., “is this regex valid?”). Importantly, these patterns reflect not just style but intent. For example, “what does” and “explain” queries almost always mapped to command interpretation tasks, while “make this better” appeared almost exclusively in text editing.
Query phrasing acts as a proxy for analyst reasoning goals, reflecting their underlying intentions and priorities, as well as their problem-solving approach and intended outcomes..
4.2.4 Alignment with Professional Cybersecurity Frameworks
Building on the task and query subject, we explored whether analyst queries aligned with established cybersecurity roles and competencies. We reviewed all queries against the NICE Framework for Cybersecurity [^3] https://www.nist.gov/itl/applied-cybersecurity/nice/nice-framework-resource-center by matching each task and subject to the NICE categories.
This comparison revealed that 93% of analyst queries could be reasonably associated with at least one NICE Task, Knowledge, or Skill element; the other 7% mostly included task themes like LLM Chat or Other. For instance, command interpretation queries align with K0805 (Knowledge of command-line tools and techniques), while email crafting aligns with S0610 (effective communication skills).
These findings suggest that LLM use was far from arbitrary; it clusters around competencies already codified by industry standards, such as the NICE framework, suggesting that LLMs could be tailored to support NICE-specific work roles rather than providing generic assistance.
4.3 Conversation Level Analysis
While the previous section focused on individual queries, this phase examines complete analyst-LLM conversations, each defined as a sequence of (temporally) linked queries. This broader lens allows us to examine how analysts pursue and refine their goals over multiple interaction steps. Our analysis examines (1) the distribution and structure of analyst-LLM conversations, highlighting common short and iterative interaction patterns; (2) the ways analysts structure queries and refine their outputs within multi-step sessions; and (3) the task transitions and sequences that characterise real-world analyst workflows when engaging with LLMs. This descriptive lens provides insight into how analysts employ LLMs as flexible, context-sensitive tools for reasoning, drafting, and refinement in SOC operations.
Our thematic coding centred on analyst queries, aligning with our objective of understanding how SOC analysts engage with LLMs in practice. Although we reviewed LLM responses to contextualise conversation-level analyst intent, we did not formally code them. This decision was based on the following considerations: (1) LLM outputs can vary in content or structure even when prompted identically, depending on deployment settings, making systematic analysis less reliable; (2) our primary research objective was to analyse analyst behaviour rather than audit or benchmark the LLM itself; and (3) defining response utility is highly analyst- and context-dependent and often lacks objective ground truth. This analytic focus allowed us to prioritise the human perspective, capturing analyst reasoning patterns and interaction strategies, without over-attributing meaning to the LLM’s behaviour, which may vary depending on deployment conditions/settings or model version.
4.3.1 Conversation Lengths
Table I summarises conversation length and the associated mean and median durations (elapsed time between first and last analyst message; LLM response time is not measured). Most conversations were brief: 57% involved just two analyst steps, and 75% contained only 2–3 queries. Only a small fraction (just over 4%) exceeded 10 steps. Durations increase with length, but longer conversations show much higher variance, with mean durations often far exceeding the median, suggesting that many multi-turn sessions are prolonged by idle periods or interruptions rather than sustained engagement. The following sections unpack how tasks and reasoning patterns unfold across these conversation lengths.
TABLE I: Summary of Conversation Lengths and Durations
Length # Conversations % of Total # Analysts Mean / Median Duration (minutes)
2 303 57.0% 35 15.8 / 2
3 96 18.0% 22 10.2 / 5
4 39 7.3% 18 28.9 / 13
5 27 5.1% 10 33.1 / 13
6 22 4.1% 13 56.4 / 22
7 13 2.4% 8 143.4 / 27.5
8 8 1.5% 5 145.0 / 89.5
9 7 1.3% 4 99.0 / 45.5
10 6 1.1% 5 806.2 / 146
11 3 0.6% 3 66.0 / 53
12 1 0.2% 1 21.0 / 21
13 1 0.2% 1 31.0 / 31
14 1 0.2% 1 1151.0 / 1151
19 1 0.2% 1 49.0 / 49
21 1 0.2% 1 298.0 / 298
33 1 0.2% 1 5770.0 / 5770
36 1 0.2% 1 211.0 / 211
37 1 0.2% 1 17728.0 / 17728
4.3.2 Most Common Two-Step Conversation Sequences
Analysis of the most frequent two-step task sequences reveals distinct usage patterns in analyst-LLM interactions. The most dominant 2-step sequence involves back-to-back instances of Command Understanding/Analysis, which occur in 16.2% (of all conversations). This suggests that analysts often explore multiple related commands in succession or seek clarification on different aspects of the same command context. In these conversations, analysts commonly asked variations of “What does [x] do?”, “Explain this [x]” and “Can you explain what [x] does?” about a command in both steps. Another recurring sequence is Command Understanding/Analysis followed by Summarising Command (3% instances; in the top 5 of all sequences), reflecting a typical workflow where analysts first understand a command and then summarise it, likely for reporting purposes.
Several editing-related sequences appear frequently. These include repeated instances of Editing/Rewriting Text (11% instances), and combinations indicating that analysts often revise content iteratively with a focus on tone or clarity. The third most frequent loop (4.7%) is repeated Code, Script & Regex Analysis, reflecting multiple passes of script writing or debugging. Several other two-step patterns (1-3% of conversations) include self-loops in Attack & Threat Intel, Tool/Software Understanding, Query/Search & Data Processing, Network Analysis, and Email Analysis.
The recurrence of short, structured sequences suggests that analysts are not simply issuing isolated queries. Instead, they use the LLM to fill in gaps as they piece together the context of an event. For example, repeated command understanding shows how analysts incrementally contextualise related commands, while rewriting or summarising sequences often reflect the need to refine language or tone for reporting or documentation. This behaviour reinforces the LLM’s role as a flexible tool: analysts use it to interpret telemetry or polish content in small, focused iterations that support the analysts’ larger investigative or communicative goals.
4.3.3 Three-Step Conversation Sequences
Among conversations consisting of three analyst steps, we observe multiple instances where analysts remain focused on a single task theme throughout. For example, we observed [Editing/Rewriting Text, Editing/Rewriting Text, Editing/Rewriting Text] sequences, where analysts revised text across multiple iterations (5%). These included sequential refinements such as “Make this better [x]”, followed by “Make it shorter [x]” and then “Make it sound better [x]”. Such sequences illustrate how analysts leverage the LLM as a drafting and refinement tool, shaping content, such as alerts, investigation notes, or emails, over several small adjustments.
There were triplets of command understanding-related queries (4%), typically involving repeated or slightly varied versions of prompts such as “What does [x] do?” or “Explain [x]”. Similar to earlier discussions, this reveals a pattern of iterative probing, where analysts utilise LLMs to interpret one or more commands or their components.
Interestingly, a small number of sequences involved transitions between different task themes (e.g., from command analysis to text editing, or alternating between command and tool understanding), highlighting cases where the analyst first interprets a command and then translates that interpretation/understanding into an incident narrative.
Overall, these three-step interactions reveal how analysts engage the LLM to incrementally make sense of complex information, refine their articulation, and support analytical workflows. The sequences reflect real-world cognitive processes, particularly when dealing with uncertainty or unfamiliar commands or scenarios.
4.3.4 Longer Conversations
While shorter interactions (lengths 2-3) were most common in our dataset, several conversations extended beyond three exchanges. Although individually less frequent, these longer interactions offer insight into how analysts build upon prior queries to progressively deepen understanding, integrate multiple concepts, or iteratively refine their outputs.
These longer conversations spanned a wide range of tasks. Some involved exploratory learning or upskilling, such as understanding machine learning fundamentals, understanding authentication mechanisms, or the Remote Desktop Protocol. Others reflected hands-on problem-solving, such as writing or debugging scripts, analysing remote shell behaviour, analysing registry modifications, or investigating potentially malicious behaviour (e.g., detecting potential ransomware). A distinct subset of longer conversations centred on professional development, likely related to CVs or performance reviews. These cases highlight the LLM’s use beyond core SOC functions, supporting analysts in broader professional contexts.
Many extended interactions show a progression from low-level technical understanding to higher-order reasoning. For example, conversations with four or more consecutive Command Understanding/Analysis steps suggest thorough exploration and careful reasoning of related commands in high-stakes environments where misinterpreting command behaviour could lead to erroneous threat assessments. Other sequences chain distinct but related subtasks, combinations such as File Understanding $\rightarrow$ Command Understanding/Analysis $\rightarrow$ Threat Analysis $\rightarrow$ Summarise LLM Output indicate layered workflows in which an analyst first clarifies data sources or command behaviour, then assesses threat implications, and finally requests help formulating an incident report or alert. We also observed iterative development conversations with repeated cycles of Code Writing and Understanding, Understanding Errors, and Query Writing, suggesting a debugging or prototyping mindset in which the LLM supports implementation, error checking, refinement, and formatting. Finally, several multi-step interactions centred on communication tasks, Editing/Rewriting Text, Improving Phrasing/Clarity, and Generating Empathetic Emails, combine structural edits with tone adjustment and simplification, highlighting the LLM’s utility in producing polished, context-appropriate communication for internal and external stakeholders.
Across all conversation lengths, a picture emerges of how SOC analysts engage with LLMs: as iterative, context-sensitive tools for reasoning, drafting, and refinement. Analysts often employ multiple steps to make sense of complex information and/or refine outputs. They use LLMs in deliberate, task-oriented ways, typically through short, focused exchanges that involve probing for interpretation, refining written outputs, or shifting between related tasks (from sense-making to writing).
4.4 Analyst Level Analysis
Our analysis reveals substantial variation in how individual analysts engaged with the LLM, both in terms of query volume and the thematic breadth of their interactions. As noted previously, analysts were grouped into three distinct clusters: low-usage analysts (Cluster 0, $n=34$), moderate users (Cluster 1, $n=10$), and a single high-volume outlier (Cluster 2, $n=1$; most active analyst).
We examined daily LLM query activity across all analysts (Figure 6) and observed distinct engagement patterns. First, one analyst (A68f) used the system consistently and at high volume across multiple months. Second, a group of moderate users (e.g., Ac9a, A5b3) engaged in concentrated bursts of activity, likely linked to specific investigations. In contrast, most analysts issued only a handful of queries on scattered days, suggesting sporadic or task-specific usage. These behavioural differences support cluster-based segmentation and reveal diverse styles of LLM integration within the analyst population.
Analysts varied markedly in the task domains they engaged with through the LLM. The most prolific user, A68f, demonstrated broad usage across multiple domains, particularly Command Understanding, Analysis & Generation, LLM Chat/System Queries, and Text Generation & Editing. This suggests deep LLM integration into technical investigations and communication-related tasks. In contrast, Ac9a concentrated overwhelmingly on Text Generation & Editing, with limited engagement in command analysis, indicating a primarily communicative or documentation-focused use.
A30d and A5b3, while heavily focused on Code, Script, and Regex Analysis, did explore other tasks moderately. Adf2 showed a more balanced profile, spanning Text Generation, Command Understanding, and Tool/System Understanding, reflecting the use of the LLM for both exploratory and functional support tasks. Analysts like A4df combined command and code-related tasks with moderate Query/Data Processing and File Analysis, suggesting use cases grounded in technical inspection.
Among lower-volume users, such as A180, Ac48, and A95e, task engagement was more narrowly focused, typically confined to one or two functional categories such as Text Editing or Command Understanding. Notably, the recurring appearance of Command Understanding and Text Generation across many analysts underscores their central role in LLM support within SOC workflows.
Figure 6: Number of queries per analyst (ordered by activity) and task theme (ordered by frequency). Analyst clusters are colour-coded: red for most active, green for moderate, and blue for low-usage analysts (same as Figure 3)
These task distinctions are mirrored by varying cognitive strategies at the pattern level. Figure 7 shows the top 5 task themes (left) and top 5 pattern themes (right) used by the 20 most active analysts. On the left side, usage is dominated by Command Understanding and Text Generation & Editing, though the mix varies by analyst. On the right, Text Processing, Editing, & Summarising is the most prevalent interaction pattern, but analysts also use a mix of patterns.
A68f, the most prolific user, led in Functional Understanding and Clarification Requests, with heavy use of LLM Interaction & System Queries. This pattern aligns with their diverse technical workload and diagnostic dialogue style. In contrast, Ac9a showed a documentation-focused usage profile via Text Processing, Editing & Summarising, alongside frequent clarification queries. Adf2 exhibited an iterative refinement strategy, issuing Correction & Error Fixing and Content Generation & Expansion queries; a cycle of generating, testing, and refining outputs.
Distinct pattern signatures also appear to reflect role-based tendencies. A30d and A5b3, for example, consistently used the LLM for Feasibility Checking, How-To Reasoning, and Causal Explanation, aligning with their hands-on technical work and exploratory debugging. Conversely, low-volume users such as A07f, A2f7, and Aae4 issued queries primarily related to Clarification or Functional Understanding, suggesting that for many, the LLM functioned as a reference assistant and a reasoning partner. However, it is important to note that a comprehensive understanding of analysts’ specific tasks and context is necessary to accurately interpret such trends; variability in queries may simply reflect the evolving nature of investigative work.
Figure 7: Top 5 task themes (left) and pattern themes (right) used by the 20 most active analysts.
We observed that LLMs are not used uniformly across analysts. Instead, individuals adapt the tool to their task demands, specific needs, and investigative styles, ranging from quick clarification and script debugging to collaborative drafting and iterative refinement. The LLM successfully supported this diversity without requiring extensive prompt engineering. This highlights that a well-configured, domain-aligned LLM can support a broad range of task types and prompt styles, enabling quick and effective interaction, even without deep prompting expertise.
4.4.1 Discussion on Potential Disengagement Patterns
The following analysis examines behavioural patterns among SOC analysts with lower engagement levels to understand where, why and how disengagements occur. We observed two distinct groups of low-engagement analysts.
Group 1: Low-Engagement, Early Drop-off Analysts (n = 13) These analysts interacted for only 1-2 months. Most (9 out of 13) showed no signs of frustration, implying exploratory or episodic use rather than dissatisfaction. However, three analysts exhibited potential breakdowns that coincided with their last few queries.
A2f7 repeatedly queried the model version (perhaps unmet transparency need), A5a6 received two consecutive erroneous answers, with a misclassification ending the session, and A6e9 disengaged immediately after the LLM refused a request. Another analyst (Ae46) ceased engagement after several command analysis queries, although no clear failure pattern was identified. Collectively, this group suggests that early disengagement occurred for multiple reasons: in some cases without clear frustration, and in others following specific model failures.
Group 2: Persistent but Mixed-Experience Analysts (n = 6) These analysts remained active for four to eight months. Two analysts (Ae32, Aae4) experienced apparent errors yet persisted, indicating that perceived utility can offset isolated failures. Other patterns suggest diverse usage motives: A9c2 moved past an early debugging issue; A214 relied on the LLM for repeated Sumo Logic refinements; A8bb may have engaged only when necessary; and A525 mixed non-SOC queries with SOC ones. This group suggests that analysts tolerate occasional mistakes, perhaps if using the LLM for specific tasks is beneficial.
While these patterns offer insights, we did not have direct access to analysts (due to privacy reasons) to confirm these. Future work should involve observing SOC analysts in their environment and conducting follow-up interviews to validate reasons for disengagement. This will be essential as LLM-based capabilities continue to evolve in SOCs.
4.4.2 AI-Assisted Decision-Making Perspective
Only a small proportion of analysts explicitly sought recommendations (we use recommendations here to mean classification requests, such as, “is it malicious” or “is this a threat”) from the LLM. Across the dataset, approximately 4% of all queries involved requests for such binary judgements. For example, if we only consider Command Understanding, Analysis & Generation, in total there were 950 queries (31%), and only 3% involved explicit recommendation requests, and these were issued by just seven analysts.
This suggests that most analysts do not seek recommendations from the LLM. Instead, they appear to prefer evidence or context that allows them to understand what is happening and make decisions independently.
To effectively support SOC analysts, AI systems must align with this preference for evidentiary support over prescriptive output, for example, by using machine-in-the-loop approach [39], which presents evidence for and against outcomes rather than definitive answers to support informed decisions and preserve human agency [40, 42]. However, because some analysts prefer explicit recommendations, a one-size-fits-all approach risks leaving needs unmet. Therefore, investigating the need for adaptive systems that support both interpretive and directive workflows is essential.
5 Discussion
This study set out to investigate a core question (restated briefly): How do SOC analysts utilise LLMs in their daily workflows, including the specific tasks they apply them to, how these uses align with established cybersecurity frameworks, and the conversational patterns that characterise their interactions? Below, we address this question through a multi-level analysis and conclude with a discussion of the study’s limitations and the broader implications for designing LLM-based assistance in SOCs.
5.1 Query Level: Analyst Needs and Task Focus
The progressive increase in LLM use across different tasks (Figure 6) and sustained engagement by a growing number of analysts (Figure 3) indicate that analysts incrementally integrated the LLM into their daily workflows, applying it to an increasing number of tasks (see Figure 9 in Appendix for task adoption over time). These usage patterns suggest that for some analysts the LLM delivered sufficient value to merit repeated use during investigative sessions.
The LLM fulfilled a key, hidden interpretive layer missing from current SOC tooling, particularly in interpreting low-level telemetry, such as commands, files, system processes and scripts. This highlights a key function of the LLM: acting as a real-time interpreter or explainer of low-level telemetry. During investigations, tier-1 analysts need to make sense of raw telemetry [5]. Instead of relying solely on documentation or prior knowledge, analysts used LLMs as on-demand aids to translate such artifacts into actionable context, potentially helping analysts discern whether observed activity was benign or malicious [5]. In this way, LLMs appear to reduce time-consuming lookups, ease cognitive load, and improve situational awareness.
Text processing use cases suggest a role in enhancing operational efficiency. Rather than spending time on repetitive communication or documentation tasks, analysts utilised the LLM to generate more polished and comprehensive materials for reporting and sharing. This would have offloaded non-investigative cognitive effort, supporting productivity without disrupting core investigative flow.
LLM usage is overwhelmingly concentrated in command interpretation and text processing, with most other task types showing signs of increasing adoption, but there is only preliminary evidence in our dataset to claim widespread use. Command and text processing tasks are used continuously and with growing frequency across analysts and months (see Figure 9 in the Appendix).
Query patterns reflected clear cognitive intent. Analysts’ queries were not just syntactic variations but reflected clear cognitive intent. Functional queries such as “what does [x] do” consistently mapped to interpretive tasks like command understanding, while requests like “reword this” signalled refinement goals tied to documentation. This high alignment between linguistic pattern and task type suggests that query structure can act as a proxy for the analyst’s immediate reasoning objective.
Design Implication: Surface-level telemetry is not the same as true comprehension. Embedding plain-language, just-in-time explanations into analyst workflows could close this gap. For example, embedding analyst role- and history-adaptive LLM explanations in SOC dashboards could accelerate triage, reduce cognitive load, and support evidence-based decisions. Prior research highlights the need for actionable, context-rich explanations for SOC analysts [16], and that user-tailored interfaces improve decision accuracy and satisfaction [50]. Agentic RAG [51] could be explored to enable such explanations by dynamically tailoring retrieval and generation based on the analyst’s role and recent interactions, ensuring that each explanation is contextually relevant and personalised to the analyst’s expertise and workflow.
5.2 Conversation-level: Reasoning Sequences and LLM Roles
Our analysis reveals that SOC analysts engage with LLMs through brief, focused, and task-driven conversations. The results indicate that these 2-3 step conversations are not exploratory; rather, they are concise, goal-directed exchanges grounded in real-time investigative needs. Such behaviour aligns with the notion of assisted intelligence [30], in which LLMs are used as on-demand tools for clarification, explanation, or transformation.
In operational SOCs, LLMs act as lightweight, on-demand aides rather than persistent copilots, helping analysts reach objectives without elaborate prompts. This real-world perspective contrasts with prior work focused on technical capabilities or prototypes [7, 8, 23], highlighting the need to study LLMs in live environments.
LLM use in our dataset aligns closely with established cognitive and collaborative AI role frameworks. For instance, command understanding aligns with the clarifier [32], the Assisted Intelligence role [30], and the consultant/search engine archetypes [31]. Text processing tasks correspond to the design phase in [32], and the Augmented Intelligence category [30]. Meanwhile, iterative reasoning and hypothesis testing interactions resemble the deliberative partner model [34], ExtendAI role [33] and align with Cooperative Intelligence [30]. Together, these roles underscore the varied, situational ways LLMs support cognition within the SOC environment.
Design Implication: Embed LLMs directly into SOC dashboards so analysts can summon help for specific microtasks (e.g., interpreting a suspicious command, summarising related logs, or drafting an incident note) without leaving their current workflow to reduce context switching, and suggest logical next steps (for example, offering a summary after several explanation requests).
Research Opportunity: While most analyst-LLM interactions focused on single tasks, some involved task transitions (e.g., command analysis to summarisation), raising questions about the value of task-specific LLMs and how to support seamless switching. Future studies could assess if task-specific LLMs enhance speed, transparency, and trust in SOC decision-making. Furthermore, Agentic AI could be explored to reduce context switching by assigning specialised agents to key micro-tasks within the analyst’s workflow [52], and even dynamically incorporate expert analysts [51], while carefully managing the security risks and operational complexity [53].
5.3 Analyst-Level: Usage Patterns and Integration Approaches
Our analyst-level analysis reveals key differences in LLM usage styles, cognitive strategies, and decision preferences, offering critical implications for the design, deployment, and evaluation of AI-powered SOC tooling.
LLM use varies. Analysts differed not only in the frequency of LLM engagement but also in the types of tasks they delegated and how they framed their requests. Some analysts primarily used the LLM for command interpretation and technical debugging, while others focused more narrowly on editing or summarisation tasks. These distinctions suggest an opportunity for investigating the effectiveness of role-aware, adaptive integration strategies that personalise tooling based on emerging usage patterns. Our findings extend early claims about LLMs as cognitive aides in SOC workflows [7, 8], and validate emerging theoretical framings that emphasise diverse user strategies and interaction modes [13, 35]. While prior work has focused on how LLMs can support structured information extraction or task acceleration, we demonstrate how usage patterns naturally emerge around analysts’ task orientations without the need for prompt engineering or adaptation.
LLMs function primarily as evidence interpreters, not decision-makers. Only 4% of queries sought explicit recommendations (e.g., “is this malicious?”). Instead, most analysts requested explanation, interpretation, or contextualisation, highlighting a strong preference for maintaining decision authority. This supports the design of LLMs as evaluative cognitive aides that surface relevant evidence rather than issuing prescriptive judgments. Our findings resonate with and provide support for broader concerns about over-reliance on automated suggestions [36, 37], reinforcing the need for machine-in-the-loop systems that enhance but do not override expert judgment [39, 40, 42]. We find that most analysts favour evidence interpretation over explicit recommendations. This opens doors for further research into why different analysts prefer different modes and whether there is a notable impact on their performance.
Other factors that modulate routine engagement: isolated errors and trust.
Disengagement is often episodic, not systemic. Among low-usage analysts, disengagement typically followed isolated errors or unmet expectations (e.g., refusal). Others continued using the LLM despite early failures, indicating that value outweighs occasional breakdowns. Clarifying outputs or allowing retries may reduce drop-off.
Sustained engagement hints at emerging trust. Trust underpins effective human-AI teaming, shaping user engagement over time. Prior work distinguishes situational trust (based on immediate interactions) from learned trust (built from experience) [54, 55, 56]. In our dataset, repeated use of the LLM for complex interpretive tasks, especially by some analysts, suggests the latter, where users expect reliable support in specific contexts. However, without direct feedback, we cannot confirm trust as the driving factor. Future research should combine usage data with trust metrics to clarify the nature of trust in human-AI collaboration.
Design Implication: Surface evidence, not recommendations for investigative tasks [39, 40, 33]. Because only 4% of analyst queries requested explicit malicious/benign judgements, for tasks that request such information, interfaces should default to an evidence-first mode, highlighting log excerpts, command traces, and ATT&CK linkages while omitting prescriptive labels.
Research Opportunity: Studies could investigate whether the preference for evidence-oriented outputs (4% in our study) generalise to different settings and SOC, and also compare evidence-oriented versus recommendation-oriented outputs on accuracy, response time, and calibrated trust [36, 37] to determine which support style minimises misuse yet maintains efficiency during high-pressure triage.
Answering the RQ: SOC analysts primarily use LLMs as cognitive aids to interpret technical artifacts and streamline communication using brief, goal-oriented queries that are closely aligned with the NICE Workforce Framework and MITRE ATT&CK Framework. Most analysts view the LLM as a cognitive aid rather than a decision-maker. Our findings indicate that LLMs serve as flexible, on-demand aids that enhance rather than replace analyst expertise. Analysts appear to optimise both time and LLM utility through short, high-value interactions, quickly interpreting obfuscated telemetry for situational awareness [5] and efficiently completing writing tasks, while retaining full decision-making authority.
5.4 Limitations and Future Work
This study has several limitations. First, all data were collected from a single SOC using a specific LLM deployment. SOCs vary in tools, workflows, and policies, so our findings may not generalise to other operational contexts or different LLMs. Second, we lack objective performance metrics, such as triage speed, false‐positive rates, LLM response accuracy, or overall investigation accuracy. However, it is important to note that the primary goal of this study was not to evaluate performance outcomes, but to explore how analysts interacted with the LLM in practice. High engagement levels suggest analysts found the LLM useful, but we cannot conclude whether these interactions translated into better or faster decisions. Third, there may be a novelty effect, with some analysts engaging with the LLM simply because it was new or conveniently accessible, rather than due to demonstrated utility. As such, it is difficult to distinguish genuine productivity gains from usage driven by curiosity or experimentation without further inquiry.
Future work can build on our findings in several ways. First, replicating this study across multiple SOCs with different toolchains, team sizes, and operating procedures will help assess the generalisability of observed interaction patterns and support the use of objective measures, such as time‐to‐triage. Second, integrating LLM features directly into SOC dashboards could reduce context-switching and offer deeper insight into how interface design shapes analyst-LLM interaction. Finally, longitudinal studies that track new and experienced analysts over time can reveal how trust, adoption, and productivity evolve as the novelty of the LLM wears off. By pursuing these directions, future work can move from descriptive usage patterns and provide more comprehensive insights into the role of LLMs in supporting analyst decision-making and collaboration within operational SOCs.
Ethical Consideration: Ethical approval for this study was granted by the researchers’ host organisation’s Ethics and Privacy Office (Approval #: 025/25). Analysts were informed that their interactions would be anonymously logged to understand how they engaged with the new LLM tools, and consented to this. Informed consent was obtained by the industry partner, and participation was completely voluntary. Prior to data sharing, the industry partner de-identified the data and removed sensitive information. All procedures were reviewed and approved by both the industry partner and the host organisation’s ethics and privacy teams.
6 Conclusion
This paper explored how SOC analysts interact with LLMs in an operational setting. By analysing thousands of analyst-generated queries, we found that analysts use LLMs as on-demand, task-focused cognitive aids for a variety of tasks, including explaining commands, writing scripts, or improving documentation, rather than as full-time copilots. Future tools should embed LLMs directly into SOC systems to deliver adaptive, context-sensitive assistance. From a research perspective, our work provides a grounded foundation for understanding analyst-LLM collaboration and underscores the need for outcome-based evaluations of their real-world impact on SOC decision-making.
7 Acknowledgments
This work is supported by the Commonwealth Scientific and Industrial Research Organisation (CSIRO) Collaborative Intelligence (CINTEL) Future Science Platform (FSP) and eSentire, Inc. Figures are designed using icons from Flaticon.com.